OpenStudy (karolynxx):

y=–6x+2 –12x – 2y =–4 How many solutions does this linear system have? a.one solution: (0, 0) b.one solution: (1, –4) c.no solution d.infinite number of solutions

OpenStudy (nincompoop):

so you have to solve this one variable at a time what you can do is plug in the value of y in the first equation into the second then solve \(-12 x - [ 2\times (\color{red}{-6x+2})] = -4 \)

OpenStudy (nincompoop):

do you get it?

OpenStudy (karolynxx):

i cant figure it out

OpenStudy (nincompoop):

what can't you figure out?

OpenStudy (karolynxx):

how to solve it

OpenStudy (nincompoop):

I basically showed you half of the process already

OpenStudy (karolynxx):

i put it in a graph but didnt find any point intersecting i believe its no solution

OpenStudy (nincompoop):

okay, then there is NO solution LAUGHING HYSTERICALLY INSIDE MY HEAD

pooja195 (pooja195):

y=–6x+2 <--equation 1 –12x – 2y=–4 <---equation 2 btw no solution is incorrect Plug equation 1 into equation 2 since the value of y is given -12x-2(-6x+2)=-4 solve for x

OpenStudy (karolynxx):

i solve the parantheis first?

pooja195 (pooja195):

correct

OpenStudy (karolynxx):

-18x-0=-4

pooja195 (pooja195):

No. Start by distributing \[\rm~−12x+(−2)(−6x)+(−2)(2)=−4 \]

pooja195 (pooja195):

Write the equation u have left

OpenStudy (karolynxx):

i did distribute thats how i got that

pooja195 (pooja195):

Show me.